The Woodworker has a pretax cost of debt of 6.25 percent and a return on assets of 15.5 percent. The debt-equity ratio is .41. Ignore taxes. What is the cost of equity? Please type out a step-by-by solution

A) 18.46 percent

B) 18.78 percent

C) 19.43 percent

D) 18.94 percent

E) 19.29 percent
Answer #1

WACC = Weight of Debt * Cost of Debt + Weight of Equity * Cost of Equity

Debt/Equity = 0.41

This implies, for every $1 equity, you have $0.41 of debt

Weight of debt = $0.41/($1 +$0.41) = 29.08%

Weight of Equity = 1 - 29.08% = 70.92%

15.5% = 29.08% * 6.25% + 70.92% * Cost of Equity

13.68% = 70.92% * Cost of Equity

Cost of Equity = 19.29% (Option E)